# The neutrino force at all length scales

### v1

@@ 5.2.2 $-\ell$ states
Note that $V_{0}$ from each of the three diagrams have the same sign at long distances (see Fig. 4), however there is an additional minus sign from the factor $-(2-4s_{W}^{2})$ with $V_{0}^{\rm PG}$ in Eqs. (50) and (51); this is why the penguin contribution differs a sign from self-energy and box for $\ell\geq 1$ states.
In conclusion, due to the very short range of the $Z$ force, for $\ell=1$ state, $\eta_{\nu}$ is enhanced by a factor of $\log\left(m_{Z}^{2}/(\alpha^{2}m_{e}^{2})\right)\approx 30$ compared with the $\ell=0$ state; while for $\ell\geq 2$ states, $\eta_{\nu}$ is enhanced by $\left(m_{Z}^{2}/(\alpha^{2}m_{e}^{2})\right)^{\ell-1}\gg 1$ and the neutrino force contribution is much larger than the tree-level $Z$ contribution. The result for $\ell\geq 2$ agrees with that in Ghosh:2019dmi which uses the four-Fermi approximation. One should note, however, that the current experiments are still not able to measure APV for $\ell>0$ states.
@@ 5.2.3 Accumulated charge
where $\psi_{\ell}$ is the radial wavefunction with angular momentum $\ell$. The accumulated charge characterizes the contribution to APV in the region $r<c/m_{Z}$ (where $c$ is any positive constant) normalized by the total contribution. For the tree-level $Z$ force with $s$-wave state, one obtains:
@@
=$\overset{c\gg 1}{=}1+\frac{10}{c^{2}}+{\cal O}\left(\frac{1}{c^{4}}\right).$ (72)
As opposed to the $Z$ force that is exponentially suppressed, the neutrino contribution is only quadratically suppressed at large distances. In the upper left panel of Fig. 5, we compare the accumulated charge of the $Z$ force and the neutrino forces for the $s$-wave state. The bumps of the neutrino curves correspond to the places where the neutrino forces change signs (see Fig. 4). This nontrivial feature cannot be captured by using the four-Fermi approximation of the neutrino force. In addition, we see that most of contributions of the neutrino force to the $\ell=0$ APV matrix element are located in the region $1\lesssim c\lesssim 10$. This is a very short distance compared to typical atomic scales, but is significantly larger than the range of the $Z$ force.
The accumulated charges for $-\ell$ states can be computed in a similar way. In particular, in the region $1/m_{Z}\ll r\ll 1/(\alpha m_{e})$, it scales as:
@@
The accumulated charges for the $p$-wave neutrino forces are shown in the upper right and lower left panels of Fig. 5, which grow logarithmically when $1/m_{Z}\ll r\ll 1/(\alpha m_{e})$. As $r$ exceeds the Bohr radius (corresponding to $c=m_{Z}/(\alpha m_{e})\approx 2\times 10^{7}$), all the accumulated charges tend to unit exponentially. In the lower right panel of Fig. 5, we compare the accumulated charges of $p$-wave and $d$-wave states at large distances. As one can see, for $\ell\geq 1$ states, the contribution from the neutrino force to APV matrix elements has an extensive range up to the atomic length scale.
@@ 5.3 Extension to heavy atom systems
muonium/hydrogen $:\quad\eta_{\nu}\sim\frac{g^{2}}{16\pi^{2}\left(1-4s_{W}^{2}\right)}\sim 5\%\;,$ (75)
heavy atoms $:\quad\eta_{\nu}\sim\frac{g^{2}}{16\pi^{2}}\sim 0.3\%\;.$ (76)
Remarkably, $\eta_{\nu}\sim 0.3\%$ in heavy atoms indicates that the contribution of the neutrino force is comparable to the current experimental error in APV experiments, which is about $0.35\%$ Wood:1997zq , and to the theoretical uncertainty of many-body calculations in the cesium atom, estimated to be about $0.45\%$ Porsev:2009pr ; Dzuba:2012kx .
@@ 5.4 Effects of neutrino forces on the extraction of $\sin^{2}\theta_{W}$
[…]
It is interesting to note that the deviation in the measurement of $\sin^{2}\theta_{W}$ caused by the neutrino force is the same order of magnitude when measured in muonium versus when it is measured in heavy atoms like Cesium. This is because the accidental enhancement factor $(1-4\sin^{2}\theta_{W})^{-1}$ to $\eta_{\nu}$ in muonium is cancelled by the same factor in Eq. (82), making the deviation insensitive to the true value of $\sin^{2}\theta_{W}$, just as the case of heavy atoms. Currently, the discrepancy in measurement of $\sin^{2}\theta_{W}$ between SM predicted value and the APV measurement in Cesium is about ${\cal O}(1\%)$, and in that regard $0.3\%$ is definitely not negligible. This is another indication that the contribution to APV from neutrino effects is important, even though for most of the literature it has been ignored.
@@ 6.1 Other quantum forces at one-loop level
In region (i), one can neglect the fermion mass and we have a full theory with two kinds of diagrams, namely self-energy ($\gamma$SE) and penguin ($\gamma$PG), as shown in Fig. 7. Note that in the self-energy diagram, $\chi$ can be charged-leptons or quarks; while in the penguin diagram, $\chi$ can only be the external fermion. The parity-violating forces in muonium and positronium are given by:
@@ 6.2 Contribution to APV matrix elements
which is numerically about $-2\%$ for $\chi=e$ and $-1\%$ for $\chi=\mu$.
@@ 7 Comparison with existing literature
The neutrino force contribution to APV in heavy atoms was first studied in Ref. Dzuba:2022vrv . However, only self-energy diagrams were included and the contribution from penguin diagrams were neglected (note that the box diagram is not present in the case of heavy atoms). Our result of the contribution from self-energy diagram seems to agree with that in Dzuba:2022vrv . In Ref. Dzuba:2022vrv , the authors report $\eta_{\nu}\approx-0.72\%$ in heavy atoms, but this number includes contributions from self-energy diagrams with all possible SM fermions except top quark in the loop, not just the neutrinos. By extracting pure neutrino contributions one can get $-0.17\%$, which is very close to what we obtain $\eta_{\nu}^{\rm SE}\approx-0.2\%$ in muonium and positronium systems. This indicates that the relative neutrino force contribution to APV may not change significantly with the atomic size. In this work we calculate the full-range potentials and the APV effects from neutrino penguin and box diagrams for the first time. We show that in muonium the penguin effect $\eta_{\nu}^{\rm PG}$ dominates over that of self-energy by a factor of $\eta_{\nu}^{\rm PG}/\eta_{\nu}^{\rm SE}\sim(1-4s_{W}^{2})^{-1}\sim{\cal O}(10)$. Therefore, it is reasonable that in muonium we obtain $\eta_{\nu}=\eta_{\nu}^{\rm SE}+\eta_{\nu}^{\rm PG}\approx\eta_{\nu}^{\rm PG}$. In positronium, the box diagram gives the largest contribution since it does not involve any $Z$-boson couplings that can suppress it, and the potential from box diagram $V_{0}^{\rm box}$ keeps the same sign through the whole range.
@@
Next, we compare the contribution to APV from photon penguin diagrams, which is parametrized by $\eta_{\gamma}=\eta_{\gamma}^{\text{SE}}+\eta_{\gamma}^{\text{PG}}$. The first term was calculated in Ref. Dzuba:2022vrv for heavy atoms, which found that $\eta_{\gamma}^{\text{SE}}\approx 2\%$, with the contribution from the electron $\eta_{\gamma}^{\text{SE},e}\approx 0.1\%$. Our calculation in muonium and positronium yields $\eta_{\gamma}^{\text{SE},e}\approx-2\%$. At the same time we find the total $\eta_{\gamma}^{\text{SE}}\approx-55\%$, with quarks giving a large contribution of $-52\%$, as seen in Eq. (97). The difference is attributed to the different vector couplings of the nucleus to the $Z$ boson. For heavy nucleus, the vector coupling is proportional to the weak charge $Q_{W}\approx-{\cal N}+(1-4s_{W}^{2}){\cal Z}\approx-{\cal N}$, whereas in muonium and positronium it is $1-4s_{W}^{2}$. As a result, the relative loop/tree coupling in heavy atoms is roughly ${\cal Z/(-N)}$, while that in muonium and positronium is $1/(1-4s_{W}^{2})$. This explains why the magnitude of $\eta_{\gamma}^{\rm SE}$ in muonium and positronium is about 20 times larger than that in heavy atoms. The second term $\eta_{\gamma}^{\rm PG}$ was not included in Dzuba:2022vrv and we find it only contributes approximately $-1\%$ ($-2\%$) in muonium (positronium). This is because the $\gamma{\rm PG}$ diagrams can only have charged-leptons in the loop, whose vector couplings to $Z$ boson are more suppressed than quarks. Both $\gamma{\rm SE}$ and $\gamma{\rm PG}$ diagrams have been calculated long time ago Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss as radiative corrections to the weak charge. However, the results in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss were only given at the amplitude level in the $q^{2}\to 0$ limit; the authors did not perform the Fourier transform and convolve with the atomic wave functions to get the parity-violating matrix elements.
@@
Finally, we would like to comment on the $\delta$-function approximation commonly used in the literature for calculating APV. At the amplitude level, compared with ${\cal O}(\alpha G_{F})$, the ${\cal O}(G_{F}^{2})$ contribution from neutrino loop is suppressed by $q^{2}/m_{Z}^{2}$, and can therefore be neglected in the limit $q^{2}\to 0$, as was done in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss . Taking the limit of zero momentum transfer is equivalent to setting $m_{Z}^{2}\to\infty$. This results in taking the $\delta$-function approximation: $V_{Z}\sim G_{F}m_{Z}^{2}e^{-m_{Z}r}/r\sim G_{F}\delta^{3}(r)$. This is how the radiative corrections to the weak charge are obtained Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss : By definition, the theoretical prediction of the weak charge (42) is given by the parity-violating Feynman amplitude in the $q^{2}\to 0$ limit. However, this approximation is only valid when the potential contributing to APV has a range smaller than $m_{Z}^{-1}$. In other words, the weak charge (by definition) cannot capture the long-range effect to APV at $r>m_{Z}^{-1}$. When there are long-range contributions, the correct way forward is to transform the amplitude into position space and convolve it with the atomic wavefunctions. Our calculation implies that for $s$-wave states, the generic ${\cal O}(G_{F}^{2})$ neutrino contribution to APV is actually of order $g^{2}/(16\pi^{2})$ [without including the accidental enhancement factor $1/(1-4s_{W}^{2})$] compared with the leading ${\cal O}(G_{F})$ contribution from $Z$-exchange. We thus conclude that it is the same order as the ${\cal O}(\alpha G_{F})$ terms calculated in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss . Therefore, contrary to what people thought before, the ${\cal O}(G_{F}^{2})$ neutrino contribution to APV at the one-loop level cannot be neglected.
@@ 8 Conclusions
This result underscores the importance of taking the limits in the correct order. Indeed, for distances much larger than $1/m_{Z}$, the neutrino force is negligible compared to other one-loop effects, such as the photon penguin, with a ratio of order $G_{F}^{2}/r^{5}$ versus $G_{F}\alpha/r^{3}$. However, for $r\sim 1/m_{Z}$, the full electroweak theory must be used, and the ratio between these diagrams is of the order of the coupling constants. Since the relevant scale for calculating APV for $s$-wave states is approximately $1/m_{Z}$, the neutrino corrections become as significant as other one-loop diagrams.
@@ Acknowledgements.
MG acknowledges support from the US Department of Energy grant DE-SC0010102. YG is supported in part by the NSF grant PHY- 2309456.
@@ A.1 The self-energy diagram
Notice that the minus sign from the neutrino loop has been cancelled by the minus sign from the anti-commutativity between fermion and anti-fermion fields. The one-loop integral can be worked out analytically in terms of the Passarino-Veltman functions Passarino:1978jh :
@@
×$\times\left(q_{\mu}q_{\nu}\left(2(\xi_{Z}-1)\left(2m_{Z}^{2}\xi_{Z}-(\xi_{Z}+1)q^{2}\right)B_{00}\left(q^{2};0,0\right)\right.\right.$
@@
where $x\equiv m_{Z}r$ and ${\rm Ei}(x)\equiv-\int_{-x}^{\infty}{\rm d}t\ t^{-1}e^{-t}$ is the exponential integral function.

### v2

@@ 5.2.2 Higher-$\ell$ states
Note that $V_{0}$ from each of the three diagrams have the same sign at long distances (see Fig. 4), however there is an additional minus sign from the factor $(2-4s_{W}^{2})$ with $V_{0}^{\rm PG}$ in Eqs. (51) and (52); this is why the penguin contribution differs a sign from self-energy and box for $\ell\geq 1$ states.
In conclusion, due to the very short range of the $Z$ force, for $\ell=1$ state, $\eta_{\nu}$ is enhanced by a factor of $\log\left(m_{Z}^{2}/(\alpha^{2}m_{e}^{2})\right)\approx 30$ compared with the $\ell=0$ state; while for $\ell\geq 2$ states, $\eta_{\nu}$ is enhanced by $\left(m_{Z}^{2}/(\alpha^{2}m_{e}^{2})\right)^{\ell-1}\gg 1$ and the neutrino force contribution is much larger than the tree-level $Z$ contribution. The result for $\ell\geq 2$ agrees with that in Ghosh:2019dmi which uses the four-Fermi approximation.
One should note, however, that the current experiments are still not able to measure APV for $\ell>0$ states. Moreover, for heavy atoms, the relativistic corrections and electron correlations are significant, which enhances the wavefunctions around the origin, and thus the results in Eqs. (67)-(70) might be changed.
@@ 5.2.3 Accumulated charge
where $\psi_{\ell}$ is the radial wavefunction with angular momentum $\ell$. The accumulated charge characterizes the contribution to APV in the region $r<c/m_{Z}$ (where $c$ is any positive constant) normalized by the total contribution. For the tree-level $Z$ force with $s$-states, one obtains:
@@
$\overset{c\gg 1}{=}1+\frac{10}{c^{2}}+{\cal O}\left(\frac{1}{c^{4}}\right).$ (73)
As opposed to the $Z$ force that is exponentially suppressed, the neutrino contribution is only quadratically suppressed at large distances. In the upper left panel of Fig. 5, we compare the accumulated charge of the $Z$ force and the neutrino forces for the $s$-state. The bumps of the neutrino curves correspond to the places where the neutrino forces change signs (see Fig. 4). This nontrivial feature cannot be captured by using the four-Fermi approximation of the neutrino force. In addition, we see that most of contributions of the neutrino force to the $\ell=0$ APV matrix element are located in the region $1\lesssim c\lesssim 10$. This is a very short distance compared to typical atomic scales, but is significantly larger than the range of the $Z$ force.
The accumulated charges for higher-$\ell$ states can be computed in a similar way. In particular, in the region $1/m_{Z}\ll r\ll 1/(\alpha m_{e})$, it scales as:
@@
The accumulated charges for the $p$-wave neutrino forces are shown in the upper right and lower left panels of Fig. 5, which grow logarithmically when $1/m_{Z}\ll r\ll 1/(\alpha m_{e})$. As $r$ exceeds the Bohr radius (corresponding to $c=m_{Z}/(\alpha m_{e})\approx 2\times 10^{7}$), all the accumulated charges tend to unit exponentially. In the lower right panel of Fig. 5, we compare the accumulated charges of the $p$-state and the $d$-state at large distances. As one can see, for $\ell\geq 1$ states, the contribution from the neutrino force to APV matrix elements has an extensive range up to the atomic length scale.
@@ 5.3 Extension to heavy atom systems
muonium/hydrogen $:\quad\left|\eta_{\nu}\right|\sim\frac{g^{2}}{16\pi^{2}\left(1-4s_{W}^{2}\right)}\sim 5\%\;,$ (76)
heavy atoms $:\quad\left|\eta_{\nu}\right|\sim\frac{g^{2}}{16\pi^{2}}\sim 0.3\%\;.$ (77)
Remarkably, $\eta_{\nu}\sim 0.3\%$ in heavy atoms indicates that the contribution of the neutrino force is comparable to the current experimental error in APV experiments, which is about $0.35\%$ Wood:1997zq , and to the theoretical uncertainty of many-body calculations in the cesium atom, estimated to be about $0.45\%$ Porsev:2009pr ; Dzuba:2012kx . Note that this $\sim 0.3\%$ effect refers to the contribution from two-neutrino exchange only, and there could be other comparable contributions to APV at the one-loop level that are not discussed here. Yet, we emphasize that the neutrino force contribution we computed above is nontrivial, in the sense that it cannot be absorbed into the tree-level vertex, is independent of the renormalization scheme, and was missed in the previous literature Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss .
@@ 5.4 Effects of neutrino forces on the extraction of $\sin^{2}\theta_{W}$
heavy atoms $:\quad\left({\sin^{2}\theta_{W}^{\prime}-\sin^{2}\theta_{W}\over\sin^{2}\theta_{W}}\right)_{\nu}\approx\frac{3{\cal N}}{{\cal N}+3{\cal Z}}\left(-\frac{g^{2}}{16\pi^{2}}\right)\overset{\text{${}^{133}$Cs}}{\approx}-0.3\%\;.$ (87)
It is interesting to note that the deviation in the measurement of $\sin^{2}\theta_{W}$ caused by the neutrino force is the same order of magnitude when measured in muonium versus when it is measured in heavy atoms like Cesium. This is because the accidental enhancement factor $(1-4\sin^{2}\theta_{W})^{-1}$ to $\eta_{\nu}$ in muonium is canceled by the same factor in Eq. (83), making the deviation insensitive to the true value of $\sin^{2}\theta_{W}$, just as the case of heavy atoms. Currently, the discrepancy in measurement of $\sin^{2}\theta_{W}$ between SM predicted value and the APV measurement in Cesium is about ${\cal O}(1\%)$, and in that regard $0.3\%$ is definitely not negligible. This is another indication that the contribution to APV from neutrino effects is important, even though for most of the literature it has been ignored.
@@ 6.1 Other quantum forces at one-loop level
[…]
@@ 6.2 Contribution to APV matrix elements
which is numerically about $2\%$ for $\chi=e$ and $1\%$ for $\chi=\mu$.
@@ 7 Comparison with existing literature
The neutrino force contribution to APV in heavy atoms was first studied in Ref. Dzuba:2022vrv . However, only self-energy diagrams were included and the contribution from penguin diagrams were neglected (note that the box diagram is not present in the case of heavy atoms). Our result of the contribution from self-energy diagram seems to agree with that in Dzuba:2022vrv . In Ref. Dzuba:2022vrv , the authors report $\eta_{\nu}\approx-0.72\%$ in heavy atoms, but this number includes contributions from self-energy diagrams with all possible SM fermions except top quark in the loop, not just the neutrinos. By extracting pure neutrino contributions one can get $0.17\%$, which is very close to what we obtain $\eta_{\nu}^{\rm SE}\approx-0.2\%$ in muonium and positronium systems. This indicates that the relative neutrino force contribution to APV may not change significantly with the atomic size. In this work we calculate the full-range potentials and the APV effects from neutrino penguin and box diagrams for the first time. We show that in muonium the penguin effect $\eta_{\nu}^{\rm PG}$ dominates over that of self-energy by a factor of $\eta_{\nu}^{\rm PG}/\eta_{\nu}^{\rm SE}\sim(1-4s_{W}^{2})^{-1}\sim{\cal O}(10)$. Therefore, it is reasonable that in muonium we obtain $\eta_{\nu}=\eta_{\nu}^{\rm SE}+\eta_{\nu}^{\rm PG}\approx\eta_{\nu}^{\rm PG}$. In positronium, the box diagram gives the largest contribution since it does not involve any $Z$-boson couplings that can suppress it, and the potential from box diagram $V_{0}^{\rm box}$ keeps the same sign through the whole range.
@@
Next, we compare the contribution to APV from photon penguin diagrams, which is parametrized by $\eta_{\gamma}=\eta_{\gamma}^{\text{SE}}+\eta_{\gamma}^{\text{PG}}$. The first term was calculated in Ref. Dzuba:2022vrv for heavy atoms, which found that $\eta_{\gamma}^{\text{SE}}\approx 2\%$, with the contribution from the electron $\eta_{\gamma}^{\text{SE},e}\approx 0.1\%$. Our calculation in muonium and positronium yields $\eta_{\gamma}^{\text{SE},e}\approx-2\%$. At the same time we find the total $\eta_{\gamma}^{\text{SE}}\approx-55\%$, with quarks giving a large contribution of $52\%$, as seen in Eq. (98). The difference is attributed to the different vector couplings of the nucleus to the $Z$ boson. For heavy nucleus, the vector coupling is proportional to the weak charge $Q_{W}\approx-{\cal N}+(1-4s_{W}^{2}){\cal Z}\approx-{\cal N}$, whereas in muonium and positronium it is $1-4s_{W}^{2}$. As a result, the relative loop/tree coupling in heavy atoms is roughly $\cal Z/(-N)$}, while that in muonium and positronium is $1/(1-4s_{W}^{2})$. This explains why the magnitude of $\eta_{\gamma}^{\rm SE}$ in muonium and positronium is about 20 times larger than that in heavy atoms. The second term $\eta_{\gamma}^{\rm PG}$ was not included in Dzuba:2022vrv and we find it only contributes approximately $1\%$ ($2\%$) in muonium (positronium). This is because the $\gamma{\rm PG}$ diagrams can only have charged-leptons in the loop, whose vector couplings to $Z$ boson are more suppressed than quarks. Both $\gamma{\rm SE}$ and $\gamma{\rm PG}$ diagrams have been calculated long time ago Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss as radiative corrections to the weak charge. However, the results in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss were only given at the amplitude level in the $q^{2}\to 0$ limit; the authors did not perform the Fourier transform and convolve with the atomic wave functions to get the parity-violating matrix elements.
@@
Finally, we would like to comment on the zero-momentum transfer limit that is commonly used in the literature to calculate APV. At the one-loop amplitude level, compared with ${\cal O}(\alpha G_{F})$, the ${\cal O}(G_{F}^{2})$ contribution from two-neutrino exchange is suppressed by $q^{2}/m_{Z}^{2}$, and therefore may be naively neglected in the limit of $q^{2}\to 0$. Taking the limit of zero momentum transfer is equivalent to setting $m_{Z}^{2}\to\infty$. This results in taking the $\delta$-function approximation to the $Z$ force: $V_{Z}\sim G_{F}m_{Z}^{2}e^{-m_{Z}r}/r\sim G_{F}\delta^{3}(r)$. This is usually how the radiative corrections to the weak charge are obtained Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss : the theoretical prediction of the weak charge (43) is given by the parity-violating Feynman amplitude in the $q^{2}\to 0$ limit. However, this is a good approximation only when the atomic wavefunctions do not change rapidly around the origin and when the potential decays quickly at $r>1/m_{Z}$. In the more general case, the weak charge cannot capture the APV effect at $r>1/m_{Z}$, because it does not contain information about atomic wavefunctions away from the origin. When there exist long-range parity-violating interactions, the more general method is to transform the amplitude into position space and then convolve it with the atomic wavefunctions. As explicitly shown in Fig. 5, for the $s$-state, the range of the neutrino contribution to APV is concentrated at very short distances compared to the atomic length scale (though still longer than the $Z$-force range), and the $\delta$-function approximation still holds, while for $\ell\geq 1$ states, the neutrino contribution to APV has an extensive range across the entire atomic length scale. Quantitatively, our calculation implies that for $s$-states, the generic ${\cal O}(G_{F}^{2})$ neutrino contribution to APV is actually of order $g^{2}/(16\pi^{2})$ [without including the accidental enhancement factor $1/(1-4s_{W}^{2})$] compared with the leading ${\cal O}(G_{F})$ contribution from $Z$-exchange. We thus conclude that it is of the same order as the ${\cal O}(\alpha G_{F})$ terms calculated in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss . Therefore, the ${\cal O}(G_{F}^{2})$ neutrino contribution to APV cannot be neglected at the one-loop level.
@@ 8 Conclusions
This result underscores the importance of taking the limits in the correct order. Indeed, for distances much larger than $1/m_{Z}$, the neutrino force is negligible compared to other one-loop effects, such as the photon penguin, with a ratio of order $G_{F}^{2}/r^{5}$ versus $G_{F}\alpha/r^{3}$. However, for $r\sim 1/m_{Z}$, the full electroweak theory must be used, and the ratio between these diagrams is of the order of the coupling constants. Since the relevant scale for calculating APV for $s$-states is approximately $1/m_{Z}$, the neutrino corrections become as significant as other one-loop diagrams.
@@ Acknowledgements.
We would like to thank Zackaria Chacko, Anson Hook, Yushan Su, Raman Sundrum, and Ken Van Tilburg for helpful discussions and feedback. MG acknowledges support from the US Department of Energy grant DE-SC0010102. YG is supported in part by the NSF grant PHY- 2309456.
@@ A.1 The self-energy diagram
Notice that the minus sign from the neutrino loop has been canceled by the minus sign from the anti-commutativity between fermion and anti-fermion fields. The one-loop integral can be worked out analytically in terms of the Passarino-Veltman functions Passarino:1978jh :
@@
[…]
@@
where $x\equiv m_{Z}r$ and ${\rm Ei}(x)\equiv-\int_{-x}^{\infty}{\rm d}t\ t^{-1}e^{-t}$ is the exponential integral function. In the first line of Eq. (A.1), there is a double pole caused by the $Z$ propagators. In order to get the finite result in the second line, the denominator needs to be shifted by an infinitesimally small imaginary part, which corresponds to the prescription from the Feynman propagator (see Eq. (179)) — a careful treatment of the integral in the presence of additional poles is provided in Appendix C.